### v2

# Resolving phonon-mediated superconducting pairing symmetries from first-principles calculation

Abstract
The quest for topological superconductors triggers revived interests in resolving non-$s$-wave pairing channels mediated by phonons. While density functional theory and denstify functional perturbtaion theory have established a powerful framework to calculate electron-phonon couplings in real materials in a first-principles way, its application is largely limited to conventional $s$-wave superconductivity. Here, we formulate an efficient and simple-to-use algorithm for first-principles pairing channel analysis, and apply it to several representative material systems.
I Introduction
In the past two decades, the first-principles calculation within the framework of density functional theory (DFT) has reached a status to reliably describe not only the normal states of a wide range of materials, but also conventional superconductivity (SC) mediated by phonons [1]. Calculating the electron-phonon couplings (EPCs) from first principle is rapidly reaching maturity, thanks to the development of density functional perturbation theory (DFPT) [2].
It is commonly known that the predominating pairing channel from phonons is of a $s$-wave symmetry, and the Cooper pairs form spin singlets. In practice, the spin degrees of freedom of EPCs are thus ignored in most first-principles calculations. To predict the superconducting transition temperature (Tc), the $s$-wave channel is implicitly assumed, either in McMillan-Allen-Dynes formula [3] or the self-consistent solution of Migdal-Eliashberg equations [4].
Recently, the potential role of phonons in achieving unconventional superconductivity attracts revived theoretical attentions. The concern is that while electron-electron Coulomb repulsion always suppresses the conventional $s$-wave pairing, it can be collaborative with phonons in stabilizing other pairing symmetries, e.g. in a topological superconductor [5, 6, 7], or in high-Tc cuprate and pnictides superconductors [8, 9, 10]. In these cases, the subleading non-$s$-wave pairing channels from EPC should be taken into account.
First-principles analysis along this direction is very limited. The pioneering calculation by Wan and Savrasov [6] on doped Bi2Se3 employed a set of orthogonal polynomials defined on the Fermi surface (FS), known as FS harmonics [11], to project EPC to different pairing channels. The results clearly indicate that non-$s$-wave pairing information is readily encoded in the standard DFPT EPC data. However, the construction of FS harmonics is tedious, especially for complicated FS’s. The functions depend on not only the crystal structure, but also FS geometry, which hinders high-throughput applications of the method.
Nomoto et al. [12] extended DFT for superconductivity (SCDFT) [13, 14] to deal with odd-parity pairing in an ab initio way. The treatment is exact when $S_{z}$ is a good quantum number, e.g. without the spin-orbit coupling (SOC). When SOC is strong, however, the exact forms of odd-parity pairing in the basis of Kohn-Sham (K-S) eigenstates are not known a priori.
The primary goal of this Article is to formulate an efficient and simple-to-use algorithm for first-principles EPC pairing-channel analysis. The key is to properly solve gauge indeterminancy of the EPCs. After that, we show that the pairing channels can be resolved by standard matrix diagonalization. We call this approach the “direct diagonalization method” (DDM). In Sec. II, we will first review the general theory of phonon-mediated pairing interactions, and then discuss the gauge complexity. In the end, we present the DDM method. In Sec. III, the DDM method is applied to several realistic material systems. Section IV concludes this Article.
II Theoretical framework
II.1 General theory
We start by considering a system with inversion ($\mathcal{I}$) and time-reversal ($\mathcal{T}$) symmetries. Without $\mathcal{I}$, the definition of parity is missing, and the gap function is in general a mixture of odd and even parities. Without $\mathcal{T}$, the scenario is complicated by the coexistence of magnetism, which we defer for future works. The combination of $\mathcal{I}$ and $\mathcal{T}$ enforces that every eigenstate is at least doubly degenerate. Every K-S eigenstates can then be labeled by a band index $n$, the lattice momentum $\mathbf{k}$ and a pseudospin variable $s=\pm$.
The DFT+DFPT calculations quantify the following EPC Hamiltonian:
$H_{ep}=\sum_{n,m,\mathbf{k,q},s,s^{\prime}\upsilon}g_{n\mathbf{k}s,m\mathbf{k+q}s^{\prime}}^{\upsilon}c_{m\mathbf{k+q}s^{\prime}}^{\dagger}c_{n\mathbf{k}s}(b_{\upsilon,\mathbf{q}}^{\dagger}+b_{\upsilon,-\mathbf{q}}),$
in which $c_{n\mathbf{k}s}$ ($b_{\upsilon,q}$) is the annihilation operator of the electron (phonon) eigenstate. The phonons are labeled by a mode index $\upsilon$ and the lattice vection $\mathbf{q}$. $g_{n\mathbf{k}s,m\mathbf{k+q}s^{\prime}}^{\upsilon}$ is the EPC coefficient.
Within the second-order perturbation, the EPC generates the pairing interaction:
$H_{pair}=-\sum_{n\mathbf{k},m\mathbf{k}^{\prime},s_{1,2,3,4}}V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})$ (1)
$c_{m\mathbf{k}^{\prime}s_{3}}^{\dagger}c_{m,\mathbf{-k}^{\prime},s_{4}}^{\dagger}c_{n,\mathbf{-k},s_{2}}c_{n\mathbf{k}s_{1}},$
with
$V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})$ $=$ $\sum_{\upsilon}\frac{g_{n\mathbf{k}s_{1},m\mathbf{k^{\prime}}s_{3}}^{\upsilon}g_{n\mathbf{-k}s_{2},m\mathbf{-k^{\prime}}s_{4}}^{\upsilon}}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}$ (2)
$\times$ $\delta(\epsilon_{n\mathbf{k}}-\epsilon_{F})\delta(\epsilon_{m\mathbf{k}\textquoteright}-\epsilon_{F}),$
in which $\epsilon_{n\mathbf{k}}$ ($\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}$) is the electron (phonon) eigen-energy, and we restrict the involved electrons residing at the Fermi energy ($\epsilon_{F}$).
To bridge the interaction to a matrix analysis problem, it is convenient to view $V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})$ as the components of a matrix $\mathbb{V}$ and define a vector operator $\hat{\Delta}$ with its components
$\hat{\Delta}(n\mathbf{k}s_{1}s_{2})=c_{n,\mathbf{-k},s_{2}}c_{n\mathbf{k}s_{1}},$ (3)
which physically annihilates a Copper pair consisting of electrons at the $|n\mathbf{k}s_{1}\rangle$ and $|n,\mathbf{-k},s_{2}\rangle$ states. Its ground-state expectation value $\langle\hat{\Delta}(n\mathbf{k}s_{1}s_{2})\rangle$ reflects the SC order parameter. Eq. (1) is then transformed into a compact quadratic form:
$H_{pair}=-\hat{\Delta}^{\dagger}\mathbb{V}\hat{\Delta}.$ (4)
Roughly, diagonalizing the kernel $\mathbb{V}$ is supposed to decompose different pairing channels.
II.2 Gauge complexity
The complexity is that EPCs are routinely evaluated on a $k$-mesh in the full Brillouin zone, where K-S orbitals at each $k$ point is treated independently with “gauge” indeterminacy. It is straightforward to see that $\hat{\Delta}(n\mathbf{k}s_{1}s_{2})$ is gauge dependent by adding an arbitrary U(1) phase to $c_{n\mathbf{k}s_{1}}$. More critically, the double degeneracy allows any recombination of the degenerate K-S orbitals as the eigenstate. Therefore, an arbitrary SU(2) rotation usually also presents in the first-principles data, especially for strong SOC systems, which hinders a textbook definition of the pairing symmetry assuming $s_{1},s_{2}$ in $\hat{\Delta}$ refer to a global spin axis. It is worth mentioning that maximally localized Wannier function interpolation is a powerful procedure for ab initio gauge fixing [15], but to properly specify the pseudo spin index, especially for strong SOC systems, could still be tricky.
II.3 DDM solution
The main advantage of the DDM method is to avoid specifying the gauge condition, such that the computational procedure could label the pairing channel in any gauge. Let us denote
$\mathcal{T}c_{n\mathbf{k}s}\equiv s\tilde{c}_{n,-\mathbf{k},-s},$ (5)
in which the sign $s$ before $\tilde{c}$ is used to explicitly track the $\mathcal{T}^{2}=-1$ condition for an electron. The tilde symbol implies that $(\tilde{c}_{n,-\mathbf{k},+},\tilde{c}_{n,-\mathbf{k},-})$ is allowed to differ from $(c_{n,-\mathbf{k},+},c_{n,-\mathbf{k},-})$, i.e. the numerical K-S states at $-\mathbf{k}$ without gauge fixing, by an arbitrary U(1)$\times$×SU(2) transformation. Thus defined $\tilde{c}_{n,-\mathbf{k},-s}$ also relates to $c_{n\mathbf{k},-s}$ via spatial inversion as:
$\mathcal{I}c_{n\mathbf{k},-s}=e^{i\alpha_{n\mathbf{k}}}\tilde{c}_{n,-\mathbf{k},-s}.$ (6)
Note that theoretical analysis usually implicitly assumes $e^{i\alpha_{n\mathbf{k}}}=1$, which however does not automatically hold for the numerical K-S states. A U(1) transformation $c_{n\mathbf{k}\pm}\rightarrow e^{i\alpha_{n\mathbf{k}\pm}}c_{n\mathbf{k}\pm}$ will lead to an overall phase shift $\alpha_{n\mathbf{k}}=\alpha_{n\mathbf{k}+}+\alpha_{n\mathbf{k}-}$.
We rewrite Eqs. (1) and (2) into
$H_{pair}=-\sum_{n\mathbf{k},m\mathbf{k}^{\prime},s_{1,2,3,4}}V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})$ (7)
$c_{m\mathbf{k}^{\prime}s_{3}}^{\dagger}\tilde{c}_{m,\mathbf{-k}^{\prime},s_{4}}^{\dagger}\tilde{c}_{n,\mathbf{-k},s_{2}}c_{n\mathbf{k}s_{1}},$
and
$V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})$ $=$ $\sum_{\upsilon}s_{2}s_{4}\frac{g_{n\mathbf{k}s_{1},m\mathbf{k^{\prime}}s_{3}}^{\upsilon}(g_{n\mathbf{k},-s_{2},m\mathbf{k^{\prime}},-s_{4}}^{\upsilon})^{*}}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}$ (8)
$\times$ $\delta(\epsilon_{n\mathbf{k}}-\epsilon_{F})\Theta(\epsilon_{m\mathbf{k}\textquoteright}-\epsilon_{F}).$
The emergence of summand sign $s_{2}s_{4}$ is a direct consequence of $\mathcal{T}^{2}=-1$ [c.f. Eq. (5)]. A useful point is that the EPC between $\tilde{c}_{m,\mathbf{-k}^{\prime},s^{\prime}}^{\dagger}$ and $\tilde{c}_{n,\mathbf{-k},s}$ can be readily obtained from that between $c_{m,\mathbf{k}^{\prime},-s^{\prime}}^{\dagger}$ and $c_{n,\mathbf{k},-s}$ by applying $\mathcal{T}$, so the tilde states never need to be explicitly constructed.
Accordingly, Eq. (3) is rewritten as
$\hat{\Delta}(n\mathbf{k}s_{1}s_{2})=\tilde{c}_{n,\mathbf{-k},s_{2}}c_{n\mathbf{k}s_{1}}.$ (9)
This definition greatly facilitates numerical analysis of gap function symmetry, because its transformation under $\mathcal{I}$ is well-defined. Referring to Eq. (6), we have
$\mathcal{I}\hat{\Delta}(n\mathbf{k}\pm\pm)=-\hat{\Delta}(n\mathbf{k}\pm\pm)$ (10)
$\mathcal{I}\hat{\Delta}(n\mathbf{k}\pm\mp)=-\hat{\Delta}(n\mathbf{k}\mp\pm).$
Note that the undertermined U(1) phase $e^{i\alpha_{n\mathbf{k}}}$ cancels.
Now, we can construct one even-parity operator:
$\hat{\Delta}^{S}(n\mathbf{k})=\frac{1}{2}[\hat{\Delta}(n\mathbf{k}+-)-\hat{\Delta}(n\mathbf{k}-+)],$ (11)
and three odd-parity operators
$\hat{\Delta}^{T_{0}}(n\mathbf{k})$ $=$ $\frac{1}{2}[\hat{\Delta}(n\mathbf{k}+-)+\hat{\Delta}(n\mathbf{k}-+)]$
$\hat{\Delta}^{T_{\pm}}(n\mathbf{k})$ $=$ $\hat{\Delta}(n\mathbf{k}\pm\pm).$ (12)
Transforming into this basis, the kernel $\mathbb{V}$ is block diagonalized into the even-parity singlet sector ($\mathbb{V}^{S}$) and the the odd-parity triplet sector ($\mathbb{V}^{T}$). The matrix elements are:
$V^{S}(n\mathbf{k},m\mathbf{k}^{\prime})$ $=$ $\sum_{\upsilon s_{1}s_{3}}\frac{|g_{n\mathbf{k}s_{1},m\mathbf{k^{\prime}}s_{3}}^{\upsilon}|^{2}}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}$ (13)
$\times$ $\delta(\epsilon_{n\mathbf{k}}-\epsilon_{F})\Theta(\epsilon_{m\mathbf{k}\textquoteright}-\epsilon_{F}),$
and
$V^{T}(n\mathbf{k},m\mathbf{k}^{\prime})$ $=$ $\sum_{\upsilon}\begin{pmatrix}g_{n\mathbf{k}+,m\mathbf{k^{\prime}}+}^{\upsilon}(g_{n\mathbf{k}-,m\mathbf{k^{\prime}}-}^{\upsilon})^{*}&-\sum_{s}g_{n\mathbf{k}+,m\mathbf{k^{\prime}}s}^{\upsilon}(g_{n\mathbf{k}-,m\mathbf{k^{\prime}}s}^{\upsilon})^{*}&-g_{n\mathbf{k}+,m\mathbf{k^{\prime}}-}^{\upsilon}(g_{n\mathbf{k}-,m\mathbf{k^{\prime}}+}^{\upsilon})^{*}\\ -\sum_{s}g_{n\mathbf{k}s,m\mathbf{k^{\prime}}+}^{\upsilon}(g_{n\mathbf{k}s,m\mathbf{k^{\prime}}-}^{\upsilon})^{*}&\sum_{ss^{\prime}}ss^{\prime}|g_{n\mathbf{k}s_{1},m\mathbf{k^{\prime}}s_{3}}^{\upsilon}|^{2}&-\sum_{s}g_{n\mathbf{k}s,m\mathbf{k^{\prime}}-}^{\upsilon}(g_{n\mathbf{k}s,m\mathbf{k^{\prime}}+}^{\upsilon})^{*}\\ -g_{n\mathbf{k}-,m\mathbf{k^{\prime}}+}^{\upsilon}(g_{n\mathbf{k},+,m\mathbf{k^{\prime}}-}^{\upsilon})^{*}&-\sum_{s}g_{n\mathbf{k}-,m\mathbf{k^{\prime}}s}^{\upsilon}(g_{n\mathbf{k}+,m\mathbf{k^{\prime}}s}^{\upsilon})^{*}&g_{n\mathbf{k}-,m\mathbf{k^{\prime}}-}^{\upsilon}(g_{n\mathbf{k}+,m\mathbf{k^{\prime}}+}^{\upsilon})^{*}\end{pmatrix}$ (14)
$\times$ $\frac{\delta(\epsilon_{n\mathbf{k}}-\epsilon_{F})\Theta(\epsilon_{m\mathbf{k}\textquoteright}-\epsilon_{F})}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}.$
It is heuristic to inspect that the average of $V^{S}(n\mathbf{k},m\mathbf{k}^{\prime})$ normalized by the Fermi surface density of states is the isotropic dimensionless EPC strength $\lambda$ commonly used to predict the $s$-wave Tc [1], which implicitly assumes that the matrix components do not contain any singular variation. The weak-SOC approximation employed by Nomoto et al. [12] to evaluate odd-parity pairing can be viewed as neglecting the off-diagonal matrix elements in Eq. (14).
Without introducing these approximations, we proceed by directly diagonalizing $\mathbb{V}^{S(T)}$:
$\mathbb{V}^{S}|\sigma_{i}\rangle$ $=$ $\lambda^{S}_{i}|\sigma_{i}\rangle,$
$\mathbb{V}^{T}|\tau_{i}\rangle$ $=$ $\lambda^{T}_{i}|\tau_{i}\rangle.$ (15)
In practice, we non-dimensionlize $\mathbb{V}$ by $V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})\rightarrow V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})N_{n\mathbf{k}}$, where $N_{n\mathbf{k}}$ is the contribution of the $|n\mathbf{k}s\rangle$ state to the total density of states. This makes the largest $\lambda^{S}_{i}$, which we denote as $\lambda^{S}_{max}$, directly comparable to the isotropic dimensionless EPC strength $\lambda$ reported in literature. The largest $\lambda^{T}_{i}$ ($\lambda^{T}_{max}$) gives the leading odd-parity spin-triplet pairing strength.
The eigenvectors:
$|\sigma_{i}\rangle$ $=$ $[\xi_{i}(n_{1}\mathbf{k}_{1}),\xi_{i}(n_{1}\mathbf{k}_{2}),...]^{T},$ (16)
$|\tau_{i}\rangle$ $=$ $[\xi_{i}(T_{0},n_{1}\mathbf{k}_{1}),\xi_{i}(T_{+},n_{1}\mathbf{k}_{1}),\xi_{i}(T_{-},n_{1}\mathbf{k}_{1}),...]^{T}$
reflect the $\mathbf{k}$-space symmetry of the corresponding pairing channel, which can be classified by the irreducible representations (irreps.) of the symmetry group.
There is still one remaining issue. So far, $\hat{\Delta}(n\mathbf{k}s_{1}s_{2})=\tilde{c}_{n,\mathbf{-k},s_{2}}c_{n\mathbf{k}s_{1}}$ and $\hat{\Delta}(n,-\mathbf{k}s_{2}s_{1})=\tilde{c}_{n\mathbf{k}s_{1}}c_{n,\mathbf{-k},s_{2}}$ are treated as independent operators, which are actually related by an undecided gauge transformation and the fermion anticommutation relation. (Mathematically, the basis of $\mathbb{V}$ is overcomplete.) Therefore, additional constraints should be applied to the eigenvectors.
Specifically, the symmetrized even-parity operator $\hat{\Delta}^{S}(n\mathbf{k})$ is U(1) and SU(2) invariant. The constraint to $|\sigma_{i}\rangle$ is simply $\xi_{i}(n_{m}\mathbf{k}_{j})=\xi_{i}(n_{m},-\mathbf{k}_{j})$. In practice, it means that eigenvectors inconsistent with it should be discarded.
For the odd-parity sector, the eigenvector coefficients $\xi_{i}(T_{0},n_{1}\mathbf{k}_{1})$, $\xi_{i}(T_{+},n_{1}\mathbf{k}_{1})$ and $\xi_{i}(T_{-},n_{1}\mathbf{k}_{1})$ mutually transform under the SU(2) rotation like a spin-1 object. Our recipe is to find a rotation such that $\xi_{i}(T_{+},n_{1}\mathbf{k}_{1})$ and $\xi_{i}(T_{-},n_{1}\mathbf{k}_{1})$ vanish. Numerically, this is equivalent to diagonalizing a 2$\times$2 matrix: $\xi_{i}(T_{0},n_{1}\mathbf{k}_{1})\sigma_{z}+\frac{\xi_{i}(T_{+},n_{1}\mathbf{k}_{1})+\xi_{i}(T_{-},n_{1}\mathbf{k}_{1})}{2}\sigma_{x}+\frac{\xi_{i}(T_{+},n_{1}\mathbf{k}_{1})-\xi_{i}(T_{-},n_{1}\mathbf{k}_{1})}{2}\sigma_{y}$ into $\tilde{\xi}_{i}(T_{0},n_{1}\mathbf{k}_{1})\sigma_{z}$. Note that the $\mathcal{T}$ symmetry enforces $\xi_{i}(T_{+},n_{1}\mathbf{k}_{1})=\xi_{i}^{*}(T_{-},n_{1}\mathbf{k}_{1})$, so this matrix is Hermitian and $\tilde{\xi}_{i}(T_{0},n_{1}\mathbf{k}_{1})\sigma_{z}$ is real. The constraint to $|\tau_{i}\rangle$ is $\tilde{\xi}_{i}(n_{m}\mathbf{k}_{j})=-e^{i\Theta_{n_{m}\mathbf{k}_{j}}}\tilde{\xi}_{i}(n_{m},-\mathbf{k}_{j})$, in which $\Theta_{n_{m}\mathbf{k}_{j}}$ is the angle between $\langle n_{m}\mathbf{k}_{j}+|\vec{S}|n_{m}\mathbf{k}_{j}+\rangle$ and $\langle n_{m},-\mathbf{k}_{j},+|\vec{S}|n_{m},-\mathbf{k}_{j},+\rangle$ after the same rotations deciding $\tilde{\xi}_{i}(T_{0},n_{m}\mathbf{k}_{j})$ and $\tilde{\xi}_{i}(T_{0},n_{m},-\mathbf{k}_{j})$. The $\mathcal{T}$ symmetry enforces $\Theta_{n_{m}\mathbf{k}_{j}}=\pm\pi$.
Table.I Material-dependent parameters for the DFT and DFPT calculations.
Materials Energy cutoff (Ha) QE $k$-mesh QE $q$-mesh EPW $k$-mesh EPW $q$-mesh
Bi2Se3 (slab) 60 30$\times$30$\times$1 4$\times$4$\times$1 180$\times$180$\times$1 180$\times$180$\times$1
Bi2Se3 40 12$\times$12$\times$12 3$\times$3$\times$3 20$\times$20$\times$20 20$\times$20$\times$20
SnTe 60 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Bi2Pd 60 12$\times$12$\times$12 4$\times$4$\times$4 20$\times$20$\times$20 20$\times$20$\times$20
Al 80 10$\times$10$\times$10 10$\times$10$\times$10 30$\times$30$\times$30 30$\times$30$\times$30
Pd 60 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Cd 60 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Ta 60 10$\times$10$\times$10 10$\times$10$\times$10 20$\times$20$\times$20 20$\times$20$\times$20
Pb 60 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Hg 70 18$\times$18$\times$18 3$\times$3$\times$3 18$\times$18$\times$18 18$\times$18$\times$18
III Applications
Below we apply DDM to several representative materials. The input data to construct $\mathbb{V}$ are obtained from Quantum Espresso with GGA-PBE exchange correlation functional [16, 17, 18], optimized norm-conserving Vanderbilt pseudopotential [19, 20] and a 0.01 Ry Methfessel-Paxton smearing [21] for all materials. The convergence criteria are $10^{-8}(a,u)$ for lattice and atomic relaxations, and $10^{-12}(a,u)$ for electronic iterations. The EPCs from Quantum Espresso are further Wannier interpolated by using the EPW code [22]. Material-dependent setups are summarized in Tab. I. Our DDM code to analyze $\mathbb{V}$ is accessible at: [https://github.com/ZimengZeng-THU/Resolving-different-superconducting-pairing].
III.1 Bi2Se3 slab
Figure 1: (a) Atomic structure of a two-quintuple layer Bi2Se3 slab employed for the DFT+DFPT calculation; (b) Electron band structure; the shadowed region indicates the range of the electron-doped Fermi level for the EPC and DDM calculations; (c) Leading pairing strength in the spin-singlet and spin-triplet sectors; the insets schematically decompose the Fermi surfaces into two circles localized on the two surfaces of the slabs. The symmetry of the corresponding DDM eigenvector [c.f. Eq. (16)] is reflected by plotting the signs of its components at a selected pair of ±$\pm\mathbf{k}$ points.
We first consider an intuitively transparent example - the well-known 3D toplogical insulator Bi2Se3 [23]. By cleaving a slab [Fig. 1(a)], helical bands form on both surfaces [Fig. 1(b)]. Within the bulk gap, the helical surface states lead to a doubly-degenerate FS. Since the other bulk bands are fully gaped, the band index $n$ can be omitted. According to our definitions [Eqs. (5) and (9)], $\hat{\Delta}(\mathbf{k}\pm\mp)$ refers to intra-surface pairing, and $\hat{\Delta}(\mathbf{k}\pm\pm)$ refers to inter-surface pairing.
When the slab is infinitely thick, the two surfaces are decoupled. Therefore, $\langle\hat{\Delta}(\mathbf{k}\pm\pm)\rangle=0$, and there is no constraint on the relative phase of $\langle\hat{\Delta}(\mathbf{k}+-)\rangle$ and $\langle\hat{\Delta}(\mathbf{k}-+)\rangle$, indicating degeneracy between the leading $\langle\hat{\Delta}^{S}\rangle$ and $\langle\hat{\Delta}^{T_{0}}\rangle$. For a finite thick slab, Cooper pairs related by spin-conserving inter-surface scatterings, e.g. $\tilde{c}_{\mathbf{-k},+}c_{\mathbf{k},-}\rightarrow\tilde{c}_{\mathbf{k},+}c_{\mathbf{-k},-}$, tend to establish a unified phase to take advantage of the attraction, making $\Delta^{S}$ the dominating pairing channel. This argument suggests that the spatial separation of the surface states makes the the leading pairing channels from the odd-parity and even-parity sectors close in strength.
Figure  1(c) plots the DDM-predicted largest eigenvalues $\lambda_{max}^{S(T)}$ of $\mathbb{V}^{S(T)}$ in a two-quintuple layer Bi2Se3 slab as a function of the Fermi level. In terms of the irreducible representations of the slab point group (D3d), the associated eigenstates are assigned to A1g and A1u, respectively. The ratio $\lambda_{max}^{T}/\lambda_{max}^{S}$ reaches up to 76.3%, which pedagogically shows the relevance of EPC calculations for a complete understanding of non-$s$-wave superconductivity.
We note that pairing requires nonvanishing density of states, so in our calculation the Fermi level is purposely elevated away from the Dirac point [Fig. 1(b)]. The range of Fermi level for EPC calculations as shadowed in Fig. 1(b) corresponds to doping 0.043$\sim$0.052 electrons per unit cell, which however is still not enough to support significantly large intrinsic $\lambda$’s. Experimentally, proximity to a bulk superconductor is typically required to achieve surface SC [24].
III.2 Doped Bi2Se3 and SnTe
Figure 2: DDM analysis results of electron-doped Bi2Se3. The eigenvalues ($\lambda_{i}$) and eigenvectors are classified by the irreps. of the D3d point group. The color contour shows the sign and magnitude of the eigenvector components on the Fermi surface.
Figure 3: DDM analysis results of electron-doped SeTe. The eigenvalues ($\lambda_{i}$) and eigenvectors are classified by the irreps. of the Oh point group. The color contour shows the sign and magnitude of the eigenvector components on the Fermi surface.
The possible topological SC in CuxBi2Se3 [25, 26] and Sn1-xInxTe [27] has motivated several previous EPC calculations on bulk Bi2Se3 and SnTe with charge doping [28, 6, 12]. Using DDM, we can obtain a systematic differentiation of pairing channels (Figs. 2 and 3). We choose to add 0.12 (0.1) electrons per unit cell to Bi2Se3 (SnTe) in line with the previous calculations to make quantitative comparisons.
For doped Bi2Se3, the leading even-parity pairing $\lambda_{A_{1g}}$ is calculated to be 0.315, in good agreement with the isotropic $\lambda$ reported in [28]. The leading odd-parity pairing $\lambda_{Eu}$ is one order of magnitude smaller than $\lambda_{A_{1g}}$, in agreement with Namoto et al.’s qualitative conclusion based on the weak-SOC approximation [12]. We should mention that Wan and Savarasov [6] in contrast predicted large odd-parity pairings $\lambda_{A_{2u}}$ and $\lambda_{E_{u}}$ based on the FS harmonics projection. This prediction was rationalized by a singular EPC associated with the acoustic phonon at $q$=(0,0,0.04) (in terms of the reciprocal vectors). However, this singular EPC was not reproduced in either [28] or [12]. We note that our phonon linewidth calculation gives very similar results to those reported in [28] and [12]. In particular, no singularity is observed at $q$=(0,0,0.04).
With respect to doped SnTe, DDM predicts that there is no odd-parity pairing channel of the same order of magnitude in strength as the leading even-parity pairing channel. The ratio between $\lambda_{A_{1u}/E_{u}}$ and $\lambda_{A_{1g}}$ is roughly around 0.1, in agreement with [12]. The leading odd-parity gap function shown in [12] can also be nicely compared with our Eu(1) eigenstate as plotted in Fig. 3. The other odd-pairing eigenstates of comparative strength have not been shown before.
III.3 Elemental metals
Figure 4: leading even-parity(black),leading odd-parity(blue),sub-leading even-parity(red) superconducting pairing channels are calculated by our method
Figure 5: DDM analysis results of Pb. The eigenvalues ($\lambda_{i}$) and eigenvectors are classified by the irreps. of the Oh point group. The color contour shows the sign and magnitude of the eigenvector components on the Fermi surface.
Figure 4 summarizes $\lambda^{S(T)}_{max}$ of superconducting elemental metals, including Pb,Ta,Al,Pd,Hg,Cd. For all these materials, $\lambda^{S}_{max}$ obtained from the DDM method can be well compared with the isotropic dimensionless $\lambda$ reported in literature [29, 30, 31]. The ratio $\lambda_{max}^{T}/\lambda_{max}^{S}$ varies between 0.1 and 0.5. It is interesting to note that the highest ratio occurs in Pb and Hg, in which EPC is known to be quite large. In experiment, Pb is a frequently used component to search for topological superconductivity [32, 33, 34, 35]. In Fig. 5, we plot the leading pairing channels of Pb resolved by DDM. It is worth further investigations to identify the key factors for a high $\lambda_{max}^{T}/\lambda_{max}^{S}$ ratio.
IV Conclusion
In summary, we present an efficient and simple-to-use algorithm for first-principles EPC pairing-channel analysis. It is theoretically known that the leading pairing channel by EPC is $s$-wave [7]. All our first-principles results support this perception. Meanwhile, the calculations also indicate that EPC contains non-negligible non-$s$-wave channels. These quantifications and further applications in a high-throughput manner are expected to provide useful guide to realize exotic superconductivity.